# Temporal Fair Division of Indivisible Goods with Scheduling

### v1

@@ 1 Introduction
Fair division is a fundamental problem in multi-agent systems and economics, traditionally focusing on the static allocation of resources, such as splitting a cake or distributing a set of goods at a single point in time [26, 10, 17, 11, 28, 6, 12, 5]. However, many real-world allocation problems are dynamic. From cloud computing resources and food bank donations to organ exchange networks, goods often arrive over time and must be allocated online [1, 16, 8, 24, 7]. Although this approach captures the dynamic feature for real-world problems, it ignores the fairness in intermediate allocations. Therefore, the goal of temporal fair division is to maintain fairness cumulatively over the long run, rather than just instantaneously in each round.
The transition from dynamic to temporal settings introduces significant theoretical challenges. Unlike dynamic fair division, where intermediate allocations do not have to achieve fairness, temporal mechanisms must take this into consideration. Recent work has painted a bleak picture for this domain: temporal fair allocations are often impossible to guarantee without compromising efficiency or social welfare [19, 14, 13]. For instance, strong fairness concepts like Envy-Freeness up to any good (EFX) are frequently unattainable in temporal settings because an agent’s perception of fairness fluctuates wildly as new goods arrive.
These impossibility results from the temporal fair division model raise a natural question: Are these failures inherent to the problem, or are they artifacts of the model’s rigidity? Standard temporal fair division assumes that goods must be consumed or allocated immediately upon arrival. However, in many applications, such as job scheduling on a server or assigning shifts, a natural flexibility exists: A system administrator can delay a job slightly; a manager can reshuffle next week’s roster; a student affairs office can assign part-time jobs to students within a flexible interval [29, 25, 21].
@@ 1.1 Main Results
  • Allocations with Scheduling: We show that scheduling improves fairness achievability in specific cases—guaranteeing TEF1 with a buffer of at least $n/2$ rounds and enabling exact TEFX for two agents with identical days. But we prove that scheduling alone is insufficient to guarantee the general existence of either TEFX or TMMS allocations.
Due to space limitations, we defer some algorithms and proofs to the supplementary materials.
@@
TEF1 TEFX $\alpha\text{-}TEFX$ TMMS
General Settings ✗for $n\geq 3$ ✓for $n=2$ [19] ✓for $T=2$ [14] [14] $\frac{\frac{1}{2}\cdot min_{g\in O}v_{i}(g)}{min_{g\in O}v_{i}(g)+\frac{1}{2}\cdot max_{g\in O}v_{i}(g)}$ ($v_{i}(g)>0$) (Theorem 11) (Theorem 20)
Identical Days ✓for $T=3$ House Allocation Settings (Theorem 5) ✓for $n=2$ (Theorem 17) ✗for $T\geq 3$ (Theorem 6) ?for $T=2$ $\frac{1}{2}$ for $n=2$ (Theorem 12) (Theorem 16)
Generalized Binary Valuation [14] ✓for $n=2$ (Theorem 7) ✗for $n=3$ (Theorem 8) $\frac{1}{2}$ (Theorem 13) [14]
Identical Valuation [14] [14] $\frac{min_{g\in O}v(g)}{max_{g\in O}v(g)+min_{g\in O}v(g)}$ ($v(g)\in R_{\{0\}\cup[1,\infty)}$) (Theorem 14) (Theorem 20)
Bi-valued Goods ? [14] $\frac{a}{b}$ $b\geq a$ (Theorem 15) (Theorem 20)
@@
General Settings ✓* $r\geq\frac{T}{2}$ [14] (Theorem 18) (Theorem 20)
Identical Days $r\geq\frac{n}{2}$ (Theorem 17) $r\geq\lfloor\frac{T}{2}\rfloor$ and $n=2$ (Theorem 19) $r\geq\lfloor\frac{T}{2}\rfloor$ and $n=2$ (Theorem 19)
Generalized Binary Valuation ✓* [14] ✓* for $n=2$ (Theorem 7) ✓* [14]
Identical Valuation ✓* [14] (Theorem 18) (Theorem 20)
@@ 1.2 Related Works
Our work is related to online fair division [1, 7, 2, 31], where some (possibly zero) items arrive in consecutive rounds. We study a stronger version of it, known as temporal fair division, in which fairness must hold at every round [19, 14, 13]. [19] showed that maintaining EF1 for goods is impossible in uninformed setting without reallocating past items, and proposed a reallocation-based algorithm for two agents. However, they proved that and provided a polynomial-time algorithm, where at each round, the allocation is EF1 in the informed setting with two agents. [14] formalized temporal fairness for goods and chores, proving that TEFX allocations may not exist and providing TEF1 algorithms for two agents with chores or restricted settings: two item types, generalized binary valuations, and unimodal preferences. Finally, [13] explored SD-EF1, EF1, and PROP1, focusing on achieving both temporal and daily fairness. They demonstrated that, while per-round SD-EF1 is generally impossible, under the identical days setting, it is always attainable, guaranteeing a final SD-PROP1 allocation.
@@ 2 Preliminaries
An allocation $A=(A_{1},\dots,A_{n})$ is defined as an ordered partial partition of $O$, where $A_{i}$ is the goods allocated to agent $i$, such that $\forall i,j\in[N]$ where $i\neq j$, $A_{i}\cap A_{j}=\emptyset$ and $\bigcup_{i\in\mathcal{N}}A_{i}=O$. In addition, under the online fair division setting, denote $A^{t}=(A^{t}_{1},\dots,A^{t}_{n})$ as the allocation from round $1$ to after round $t$, that is $A^{t}_{i}=A^{t-1}_{i}\cup(A_{i}\cap O^{t})$.
We adopt standard fairness definitions from [3]. Envy-Freeness (EF) [30], which requires agents to value their own bundle most, is not always achievable with indivisible goods. Thus, we study two relaxations: EF1 [26, 10] allows envy if it is eliminated upon removing some item from the envied bundle, while the stronger EFX [11, 17] requires envy to be eliminated upon removing any item. Given that EFX existence is open for $n>3$, we also consider $\alpha$-EFX [28], which approximates EFX by requiring the ratio of an agent’s value to the envied bundle (minus any item) to be at least $\alpha$. Additionally, we consider Maximin Share (MMS) [10], which guarantees agents the value they could secure if they partitioned the goods themselves and received the worst bundle.In the temporal setting, we prefix these notions with “T” (i.e., TEF1, TEFX, $\alpha$-TEFX, TMMS) to indicate that fairness must hold cumulatively at every round [14].
Next, we introduce various settings that we will study.
@@
A temporal fair division instance is under the identical days setting iff $\forall t,t^{\prime}\in[T]$, there is a bijection $f_{t,t^{\prime}}:O_{t}\Rightarrow O_{t^{\prime}}$ s.t. $\forall i\in[n],\forall g\in O_{t},v_{i}(g)=v_{i}(f_{t,t^{\prime}}(g))$ (studied by [20]).
@@
A temporal fair division instance is under the generalized binary valuation setting iff $\forall t\in[T],\forall g\in O_{t},\forall i\in[n],v_{i}(g)\in\{0,b\}$ where $b\in\mathbb{R}^{+}$ (studied by [18]).
@@
A temporal fair division instance is under the bi-valued valuation setting iff $\forall t\in[T],\forall g\in O_{t},\forall i\in[n],v_{i}(g)\in\{b_{1},b_{2}\}$ where $b_{1},b_{2}\in\mathbb{R}^{+}$ (studied by [4]).
@@
A temporal fair division instance is under the identical valuation setting iff $\forall t\in[T],\forall g\in O_{t},\forall i,j\in[n],v_{i}(g)=v_{j}(g)$ (studied by [28]).
@@
A temporal fair division instance is under the house allocation setting iff $\forall t\in[T],|O_{t}|=n$, and $\forall i\in N,|A^{t}_{i}\backslash A^{t-1}_{i}|=1$ (studied by [15]).
@@ 3 Commonly Used Fair Division Algorithms
For completeness and to avoid ambiguity in our proofs, we will present the Envy-Cycle Elimination algorithm and the Round-Robin algorithm, along with their respective properties. We outline the high-level idea of each algorithm and provide the pseudocode in the appendix.
@@ 3.2 Envy-Cycle Elimination
Algorithm 4, the Envy-Cycle Elimination algorithm, will be referred to as ECE throughout the paper. Intuitively, ECE repeatedly selects the agent that is unenvied and allocates a good to them. it guarantees the existence of at least one unenvied agent by swapping bundles if an envy cycle occurs.
@@
ECE returns an EF1 allocation and runs in polynomial time [26].
Algorithm 5 will be referred to as MAX-ECE throughout the paper. MAX-ECE is a variant of ECE, where Theorem 1 still holds for MAX-ECE. MAX-ECE shares the same high-level idea as ECE, but instead of allocating any unallocated good, we allocate the maximum valued good to the unenvied agent. This modification ensures some nice properties regarding EFX and $\alpha$-EFX.
@@
MAX-ECE returns an EFX allocation for two agents if both agents have the same valuation profile [28].
@@
MAX-ECE returns a $\frac{1}{2}$-EFX allocation [12].

### v3

@@ 1 Introduction
Fair division is a fundamental problem in multi-agent systems and economics, traditionally focusing on the static allocation of resources, such as splitting a cake or distributing a set of goods at a single point in time (Lipton et al., 2004; Budish, 2011; Gourvès et al., 2014; Caragiannis et al., 2019; Plaut and Roughgarden, 2020; Aziz et al., 2019; Chan et al., 2019; Amanatidis et al., 2020). However, many real-world allocation problems are dynamic. From cloud computing resources and food bank donations to organ exchange networks, goods often arrive over time and must be allocated online (Aleksandrov et al., 2015; Ghodsi et al., 2013; Bertsimas et al., 2013; Lee et al., 2019; Benade et al., 2024). Although this approach captures the dynamic feature for real-world problems, it ignores the fairness in intermediate allocations. Therefore, the goal of temporal fair division is to maintain fairness cumulatively over the long run, rather than just instantaneously in each round.
The transition from dynamic to temporal settings introduces significant theoretical challenges. Unlike dynamic fair division, where intermediate allocations need not be fair, temporal mechanisms must account for this. Recent work has painted a bleak picture for this domain: temporal fair allocations are often impossible to guarantee without compromising efficiency or social welfare (He et al., 2019; Elkind et al., 2025; Cookson et al., 2025). For instance, strong fairness concepts like Envy-Freeness up to any good (EFX) are frequently unattainable in temporal settings because an agent’s perception of fairness fluctuates wildly as new goods arrive.
These impossibility results from the temporal fair division model raise a natural question: Are these failures inherent to the problem, or are they artifacts of the model’s rigidity? Standard temporal fair division assumes that goods must be consumed or allocated immediately upon arrival. However, in many applications, such as job scheduling on a server or assigning shifts, a natural flexibility exists: A system administrator can delay a job slightly; a manager can reshuffle next week’s roster; a student affairs office can assign part-time jobs to students within a flexible interval (Schwiegelshohn and Yahyapour, 2000; Li et al., 2021; Im and Moseley, 2020).
@@ 1.1 Main Results
  • Allocations with Scheduling: We show that scheduling improves fairness achievability in specific cases—guaranteeing TEF1 with a buffer of at least $n/2$ rounds with identical days. But we prove that scheduling alone is insufficient to guarantee the general existence of either TEFX or TMMS allocations.
@@
TEF1 TEFX $\alpha$-TEFX TMMS
General Settings ✗for $n\geq 3$ ✓for $n=2$ (He et al., 2019) ✓for $T=2$ (Elkind et al., 2025) (Elkind et al., 2025) $\frac{\frac{1}{2}\cdot min_{g\in O}v_{i}(g)}{min_{g\in O}v_{i}(g)+\frac{1}{2}\cdot max_{g\in O}v_{i}(g)}$ ($v_{i}(g)>0$) (Theorem 10) (Theorem 18)
Identical Days ✓for $T=3$ House Allocation Settings (Theorem 5) ✗for $T\geq 3$ (Theorem 6) ?for $T=2$ $\frac{1}{2}$ for $n=2$ (Theorem 11) (Theorem 15)
Generalized Binary Valuation (Elkind et al., 2025) ✓for $n=2$ (Theorem 7) $\frac{1}{2}$ (Theorem 12) (Elkind et al., 2025)
Identical Valuation (Elkind et al., 2025) (Elkind et al., 2025) $\frac{min_{g\in O}v(g)}{max_{g\in O}v(g)+min_{g\in O}v(g)}$ ($v(g)\in R_{\{0\}\cup[1,\infty)}$) (Theorem 13) (Theorem 18)
Bi-valued Goods ? (Elkind et al., 2025) $\frac{a}{b}$ $b\geq a$ (Theorem 14) (Theorem 18)
@@
General Settings ✓* $r\geq\frac{T}{2}$ (Elkind et al., 2025) (Theorem 17) (Theorem 18)
Identical Days $r\geq\frac{n}{2}$ (Theorem 16) ? ?
Generalized Binary Valuation ✓* (Elkind et al., 2025) ? ✓* (Elkind et al., 2025)
Identical Valuation ✓* (Elkind et al., 2025) (Theorem 17) (Theorem 18)
@@ 1.2 Related Works
Our work is related to online fair division (Aleksandrov et al., 2015; Benade et al., 2024; Aleksandrov and Walsh, 2017; Wang and Wei, 2026), where some (possibly zero) items arrive in consecutive rounds. We study a stronger version of it, known as temporal fair division, in which fairness must hold at every round (He et al., 2019; Elkind et al., 2025; Cookson et al., 2025). He et al. (2019) showed that maintaining EF1 for goods is impossible in an uninformed setting without reallocating past items, and proposed a reallocation-based algorithm for two agents. However, they provided a polynomial-time algorithm, where at each round, the allocation is EF1 in the informed setting with two agents. Elkind et al. (2025) formalized temporal fairness for goods and chores, proving that TEFX allocations may not exist and providing TEF1 algorithms for two agents with chores or restricted settings: two item types, generalized binary valuations, and unimodal preferences. Finally, Cookson et al. (2025) explored SD-EF1, EF1, and PROP1, focusing on achieving both temporal and daily fairness. They demonstrated that, while per-round SD-EF1 is generally impossible, under the identical days setting, it is always attainable, guaranteeing a final SD-PROP1 allocation.
@@ 2 Preliminaries
An allocation $A=(A_{1},\dots,A_{n})$ is defined as an ordered partial partition of $O$, where $A_{i}$ is the goods allocated to agent $i$, such that $\forall i,j\in[N]$ where $i\neq j$, $A_{i}\cap A_{j}=\emptyset$ and $\bigcup_{i\in\mathcal{N}}A_{i}=O$.
In addition, under the online fair division setting, denote $A^{t}=(A^{t}_{1},\dots,A^{t}_{n})$ as the allocation from round $1$ to after round $t$, that is $A^{t}_{i}=A^{t-1}_{i}\cup(A_{i}\cap O^{t})$.
In Section 4, we assume the scheduling buffer $r=1$. We defer the detailed definition of $r$ to Section 5.
To formally evaluate allocations within our temporal framework, we build upon the standard fairness taxonomy established in the static fair division literature (Amanatidis et al., 2023). The ideal benchmark is exact Envy-Freeness (EF) (Varian, 1974), which dictates that no agent strictly prefers another agent’s bundle to their own. However, when goods are indivisible, exact EF is notoriously impossible to guarantee in all instances.
This inherent limitation necessitates practical relaxations. The most common baseline is Envy-Freeness up to One Good (EF1) (Lipton et al., 2004; Budish, 2011), which permits envy provided it vanishes upon the hypothetical removal of the most valuable item from the envied agent’s bundle. A strictly stronger and more desirable standard is Envy-Freeness up to Any Good (EFX) (Caragiannis et al., 2019; Gourvès et al., 2014), requiring the envy to dissipate even if the least valuable item is removed. Because the existence of exact EFX remains a profound open problem for $n>3$ agents, the literature frequently relies on multiplicative approximations, denoted as $\alpha$-EFX (Plaut and Roughgarden, 2020). Under $\alpha$-EFX, an agent’s valuation of their own bundle must be at least a fraction $\alpha\in(0,1]$ of their valuation of the envied bundle after removing any single item. Alongside these envy-based metrics, we also evaluate Maximin Share (MMS) (Budish, 2011), a threshold-based guarantee ensuring that each agent receives a bundle worth at least what they could secure if they acted as the divider and received the worst resulting share.
The transition from static to dynamic environments profoundly disrupts the feasibility of these guarantees. In the temporal setting, we denote the perpetual enforcement of these properties by prefixing them with ”T” (i.e., TEF1, TEFX, $\alpha$-TEFX, TMMS), dictating that the respective fairness condition must be satisfied cumulatively at every discrete round $t\in[T]$ (Elkind et al., 2025).
This temporal requirement fundamentally alters the landscape of existence bounds. As our subsequent analysis will demonstrate, the relentless need to maintain immediate, round-by-round equity causes exact TEFX and TMMS to frequently collapse into impossibility across multiple valuation domains. Consequently, establishing robust $\alpha$-TEFX approximations serves as our primary mathematical vehicle for guaranteeing structured equity in these highly constrained temporal environments.
Definition 1 (Envy-Freeness (EF)).
An allocation $A=(A_{1},\dots,A_{n})$ is envy-free (EF) if $\forall i,j\in N,v_{i}(A_{i})\geq v_{i}(A_{j})$.
Definition 2 (Temporal Envy-Freeness (TEF)).
For any $t\in[T]$, an allocation $A^{t}=(A^{t}_{1},\dots,A^{t}_{n})$ is temporal envy-free (TEF) if $\forall t^{\prime}\leq t,A^{t^{\prime}}$ is EF.
Definition 3 (Envy-Freeness Up to One Good (EF1)).
An allocation $A=(A_{1},\dots,A_{n})$ is envy-free up to one good (EF1) if $\forall i,j\in N,\exists g\in A_{j}$ s.t. $v_{i}(A_{i})\geq v_{i}(A_{j}\backslash\{g\})$.
Definition 4 (Temporal Envy-Freeness Up to One Good (TEF1)).
For any $t\in[T]$, an allocation $A^{t}=(A^{t}_{1},\dots,A^{t}_{n})$ is temporal envy-free up to one good (TEF1) if $\forall t^{\prime}\leq t,A^{t^{\prime}}$ is EF1.
Definition 5 (Envy-Freeness Up to Any Good (EFX)).
An allocation $A=(A_{1},\dots,A_{n})$ is envy-free up to any good (EFX) if $\forall i,j\in N,\forall g\in A_{j}$ s.t. $v_{i}(A_{i})\geq v_{i}(A_{j}\backslash\{g\})$.
Definition 6 (Temporal Envy-Freeness Up to Any Good (TEFX)).
For any $t\in[T]$, an allocation $A^{t}=(A^{t}_{1},\dots,A^{t}_{n})$ is temporal envy-free up to any good (TEFX) if $\forall t^{\prime}\leq t,A^{t^{\prime}}$ is EFX.
Definition 7 (Approximately Envy-Freeness Up to Any Good ($\alpha$-EFX)).
Let $\alpha\in(0,1]$. An allocation $A=(A_{1},\dots,A_{n})$ is approximately envy-free up to any good ($\alpha$-EFX) if $\forall i,j\in N,\forall g\in A_{j}$ s.t. $v_{i}(A_{i})\geq\alpha\cdot v_{i}(A_{j}\backslash\{g\})$.
Definition 8 (Approximately Temporal Envy-Freeness Up to Any Good ($\alpha$-TEFX)).
For any $t\in[T]$, an allocation $A^{t}=(A^{t}_{1},\dots,A^{t}_{n})$ is approximately temporal envy-free up to any good ($\alpha$-TEFX) if $\forall t^{\prime}\leq t,A^{t^{\prime}}$ is $\alpha$-EFX.
Definition 9 (Maximin Share Fairness (MMS)).
Denote $M_{n}(O)$ to be the collection of all possible allocations of the goods in $O$ to $n$ agents. An allocation $A=(A_{1},\dots,A_{n})$ is maximin share fair (MMS) if $\forall i\in N$, $v_{i}(A_{i})\geq\mu_{i}^{n}(O)=\max_{B\in M_{n}(O)}\min_{S\in B}v_{i}(S)$.
Definition 10 (Temporal Maximin Share Fairness (TMMS)).
For any $t\in[T]$, an allocation $A^{t}=(A^{t}_{1},\dots,A^{t}_{n})$ is temporal maximin share fair (TMMS) if $\forall t^{\prime}\leq t,A^{t^{\prime}}$ is MMS.
Having established the foundational metrics of temporal fairness, we now define the specific operational constraints under which we will evaluate these properties. To systematically examine the boundaries of temporal fair division, we analyze several structured settings, beginning with the Identical Days environment (Igarashi et al., 2024).
Intuitively, this setting dictates that the exact same set of items—with identical valuation profiles—arrives at every discrete time step. While mathematically rigid, this structure directly models pervasive real-world allocation paradigms characterized by strict periodicity. For example, in cloud computing and network bandwidth distribution, a static pool of server resources or time-slots is continuously provisioned and allocated in recurring, identical cycles (Baruah et al., 1995).
@@
A temporal fair division instance is under the identical days setting iff $\forall t,t^{\prime}\in[T]$, there is a bijection $f_{t,t^{\prime}}:O_{t}\Rightarrow O_{t^{\prime}}$ s.t. $\forall i\in[n],\forall g\in O_{t},v_{i}(g)=v_{i}(f_{t,t^{\prime}}(g))$.
The second environment we investigate is the Generalized Binary Valuation setting (Halpern et al., 2020). Under this constraint, agents exhibit strictly dichotomous preferences: an agent values an arriving good at either zero or a fixed positive utility $b$. This mathematical restriction effectively models practical allocation scenarios characterized by strict, needs-based utility. In many real-world contexts, individuals possess highly polarized preferences—they either require a specific resource and derive a standard baseline benefit from it, or they find it entirely useless (Camacho et al., 2023). By reducing the valuation space to these stark binary extremes, we can isolate the algorithmic friction caused by competing absolute needs, without the noise of marginal-utility differences.
@@
A temporal fair division instance is under the generalized binary valuation setting iff $\forall t\in[T],\forall g\in O_{t},\forall i\in[n],v_{i}(g)\in\{0,b\}$ where $b\in\mathbb{R}^{+}$.
Building upon the dichotomy of generalized binary preferences, the third environment we investigate is the Bi-valued Valuation setting (Amanatidis et al., 2021). In this domain, the valuation space is restricted to exactly two strictly positive outcomes, $a$ and $b$. While this retains the structural simplicity of polarized preferences, it captures a more nuanced economic reality. In many practical scenarios, agents possess strong, categorical preferences for high-tier resources, yet they still derive a tangible, baseline utility from secondary or less-preferred goods, rather than viewing them as entirely worthless. This setting allows us to examine the resilience of temporal fairness guarantees when all items are universally desirable, but subject to strict quality or preference tiers.
@@
A temporal fair division instance is under the bi-valued valuation setting iff $\forall t\in[T],\forall g\in O_{t},\forall i\in[n],v_{i}(g)\in\{b_{1},b_{2}\}$ where $b_{1},b_{2}\in\mathbb{R}^{+}$.
The fourth environment we explore is the Identical Valuation setting (Plaut and Roughgarden, 2020). In this scenario, every agent assigns the exact same utility to any given item. This strict symmetry accurately models the distribution of universal commodities, standard-issue resources, or fungible assets, such as monetary grants, where subjective differences in preference are non-existent and allocations are purely driven by the objective value of the goods.
@@
A temporal fair division instance is under the identical valuation setting iff $\forall t\in[T],\forall g\in O_{t},\forall i,j\in[n],v_{i}(g)=v_{j}(g)$.
The final environment we explore is the House Allocation setting (Gan et al., 2019). In this highly structured domain, the number of arriving goods exactly matches the number of agents at each time step, and the mechanism is strictly constrained to allocate exactly one item to every agent per round. This framework perfectly models assignment problems governed by rigid quotas or capacity constraints—such as assigning daily shifts to employees or allocating individual dormitories—and has been extensively studied to pursue economic efficiency (Dai et al., 2024).
@@
A temporal fair division instance is under the house allocation setting iff $\forall t\in[T],|O_{t}|=n$, and $\forall i\in N$, each agent must be allocated exactly one good.
@@ 3 Commonly Used Fair Division Algorithms
In this section, we lay the methodological foundation for our work by detailing several prominent algorithmic frameworks in the fair division literature. Rather than merely listing these procedures, we examine their underlying mechanics—such as envy-cycle breaking, greedy sequential selection, and localized bundle swapping—to illustrate how they enforce strict equity constraints. We begin by reviewing classic offline mechanisms, namely Envy-Cycle Elimination (ECE), its strategic refinement MAX-ECE, and the Round-Robin (RR) protocol, which are celebrated for securing static fairness guarantees like EF1 and EFX. Subsequently, we transition to the dynamic setting by exploring Envy Balancing (EB), an algorithm explicitly designed to maintain temporal fairness (TEF1) across consecutive rounds. Together, these established protocols serve as the essential algorithmic building blocks for constructing and analyzing our novel sliding window solutions.
@@ 3.2 Envy-Cycle Elimination
Our first algorithmic building block is the classic Envy-Cycle Elimination (ECE) algorithm introduced by Lipton et al. (2004). The core principle of ECE is to iteratively allocate the next available good to an unenvied agent. Because the receiving agent is currently not envied by anyone, giving them one additional item ensures that any newly created envy is bounded by exactly that single item, thereby preserving the EF1 property.
However, during the allocation process, a state may arise where every agent is envied by at least one other agent. To resolve this and guarantee the existence of an unenvied agent, the algorithm maintains a directed envy-graph. In this graph, vertices represent agents, and a directed edge from $i$ to $j$ exists if and only if agent $i$ envies agent $j$ (i.e., $v_{i}(A_{i})<v_{i}(A_{j})$). If every agent is envied, every node in the graph has an in-degree of at least one, meaning the graph must contain a directed cycle (an envy-cycle).
The algorithm systematically breaks this cycle by performing a cyclic shift of bundles along the cycle’s edges: each agent receives the bundle of the agent they currently envy. This exchange strictly improves the utility of every agent involved in the cycle without violating the existing EF1 guarantees. By repeatedly identifying and eliminating these cycles, the envy-graph is eventually reduced to a Directed Acyclic Graph (DAG). Every DAG must contain at least one source node (an agent with an in-degree of zero), which perfectly corresponds to an unenvied agent ready to receive the next good.
@@
ECE returns an EF1 allocation and runs in polynomial time (Lipton et al., 2004).
Building upon this foundation, we consider a strategic refinement of the ECE framework known as MAX-ECE. The primary distinction lies in the item selection rule: rather than allocating an arbitrary available good to the unenvied agent, MAX-ECE mandates a greedy approach where the agent receives their most preferred available good (i.e., the good that maximizes their marginal utility).
This deterministic selection mechanism inherently preserves the baseline EF1 guarantee established in Theorem 1. More importantly, by ensuring that high-value items are distributed earlier in the allocation process, MAX-ECE yields strictly stronger fairness guarantees. As demonstrated in the subsequent theorems, this relatively simple modification allows the algorithm to achieve exact EFX for identical valuations, and a bounded approximation of EFX in the general case.
@@
MAX-ECE returns an EFX allocation for two agents if both agents have the same valuation profile (Plaut and Roughgarden, 2020).
@@
MAX-ECE returns a $\frac{1}{2}$-EFX allocation (Chan et al., 2019).